A Cr(s)|Cr3 (aq)||Fe3 (aq)|Fe(s) galvanic cell has a standard cell potential of 0.700 V. Calculate the Gibbs free energy change at 25 °C when 2.61 g of iron is deposited. Assume the concentrations in the cell remain at the standard state values of 1 M through the entire deposition process.

Calculate the maximum amount of work done by the cell on its surroundings.
Answer #1

Cr - 3e Cr3+

Fe3+ +3e Fe

Cell reaction; Cr + Fe3+ Cr3+ + Fe

Number of electrons transferred (n) = 3

When concentration of all species = 1 M

Then, Ecell = Eocell

= - nFEocell = - 3*96500*0.700 = - 202650 J/mole.

2.61 g Iron (2.61/56) mole Fe = 0.0466 mole Fe.

Then, gibbs energy change for 0.0466 mole Fe is

= 0.0466 *202650 = - 9445 J.

Work done on the surroundings (W) = - = 9445 J.